A random sample is obtained from a population with a mean of LaTeX: \mu μ = 20. After a treatment is administered to the individuals in the sample, the sample mean is M = 21.65 with a variance of s2 = 9 and standard deviation s = 3. Assuming that the sample consists of n = 36, use a two-tailed hypothesis with LaTeX: \alpha α = .05 to determine the value of t. Report the t value to two decimal places.
Answer #1

Solution,

Given that,

The null and alternative hypothesis is ,

H0 :  \mu = 20

Ha : \mu   \neq 20

M = 21.65

s = 3

n = 36

Test statistic = t =

= (M - \mu ) / s / \sqrt n

= (21.65 - 20) / 3 / \sqrt 36

t = 3.30